The locus of the complex number arg(z* + i√3) - π/4 is equal to?


1
Expert's answer
2021-08-25T09:11:48-0400

"\text{Given $\rm{arg }(z + i \sqrt{3}) = \frac\pi4$}\\
\text{Note that : $\rm{arg }(z + i \sqrt{3}) = \rm{arg }(x + i(y+ \sqrt{3})) = \tan^{-1} \frac{y + \sqrt3}{x} = \frac\pi4$} \\
\implies \frac{y + \sqrt3}{x} = \tan \frac\pi4 = 1 \\
\implies x = y + \sqrt3 \Leftrightarrow y = x - \sqrt3 \\
\text{$\therefore$ the locus of the given equation is } y = x - \sqrt3"